a rectangular block of dimension 2.0*1.0*0.5* weighs 200N calculate the maximum pressure exerted by the block on a horizontal floor
1
Expert's answer
2018-03-12T16:40:06-0400
We can find the maximum pressure exerted by the block on a horizontal floor from the formula:
P=F/A,
here, P is the maximum pressure exerted by the block on a horizontal floor, F is the weight of the block, A is the area of the block.
Then, we get:
P=F/A=(200 N)/(0.01 m∙0.005 m)=4∙〖10〗^6 Pa.